# appIReferences for Extended Technical Appendix

### v3

@@ A.1 Notation
  • $N$: Total heavy atoms per structure.
@@
  • =$\mathcal{R}=\{R_{1},...,R_{N_{\text{res}}}\}$: Partition of atoms into residues.
@@ A.2 Overall Training and Generation Workflow
9:Train $\text{Encoder}_{\theta}$ on $GraphDataset$ using coordinate reconstruction loss $\mathcal{L}_{\text{HNO}}$ (Eq. in Appendix E).
@@
[…]
@@
[…]
@@
37:Train $\epsilon_{\theta}$ on $DiffusionInputDataset$ using $\mathcal{L}_{\text{diffusion}}$ (Eq. 2).
@@ Appendix B Input Data Processing and Representation
  • Zero-Based Contiguous Atom Indexing: A new, zero-based, and contiguous indexing scheme ($10,1,\dots,N-$ 1) is created for all $N$ heavy atoms in the protein system. This re-indexing maps original PDB atom identifiers to a consistent integer range, essential for constructing graph inputs and feature matrices for the neural network.
  • Residue Definitions: Residues are also re-indexed contiguously (e.g., $10,\dots,N_{\text{res}}-$ 1). For each re-indexed residue, the file stores:
@@
  • Dihedral Angle Definitions (Atom Quadruplets): This is a critical component for calculating dihedral-based losses and analyses. For each residue, the file stores definitions for all applicable standard backbone angles ($\phi,\psi$) and sidechain angles ($\chi_{1},\chi_{2},\chi_{3},\chi_{4},\chi_{5}$). Each dihedral angle is defined by an ordered quadruplet of four atom indices. Crucially, these atom indices adhere to the new, zero-based, contiguous indexing scheme. For example, a $\phi$ angle for a specific residue would be defined by four specific integer indices from the $10\dots N-$ 1 range. This allows for unambiguous calculation of any dihedral angle directly from a set of $N$ atomic coordinates, whether they are ground truth or model-predicted. The definitions also account for residue types where certain $\chi$ angles are not present (e.g., Alanine has no $\chi$ angles, Glycine has no sidechain).
@@ Appendix D Latent Diffusion Model Details
This section details the Denoising Diffusion Probabilistic Model (DDPM) [47] utilized in our framework. The DDPM operates on the pooled latent embeddings $\mathbf{h}_{0}\in\mathbb{R}^{d_{p}}$ (where $d_{p}$ is the dimension of the pooled latent space, dependent on the pooling strategy) obtained from the encoder and pooling stages. Specific architectures and hyperparameters for the denoising network $\epsilon_{\theta}$ are discussed in the Extended Technical Appendix (Section I).
@@ F.1 Dataset and MD Simulation Protocol
The conformational dataset for the human Dopamine D2 receptor (D2R) was generated from all-atom Molecular Dynamics (MD) simulations. System Preparation: Simulations were initiated from the cryo-EM structure of the D2R in complex with the inverse agonist risperidone (PDB ID: 6CM4 [66]). The risperidone ligand was removed, and the third intracellular loop (ICL3), which is typically flexible or unresolved, was remodeled using RosettaRemodel [67] to represent an apo-like state. The final remodeled D2R structure consisted of 273 residues, comprising 2191 heavy atoms after selection for the simulation system; hydrogen atoms were not explicitly included as input features to our generative model, which focuses on heavy-atom representations. The D2R protein was then embedded in a 1-palmitoyl-2-oleoyl-sn-glycero-3-phosphocholine (POPC) lipid bilayer using the CHARMM-GUI Membrane Builder [68]. The system was solvated with TIP3P water [69] and neutralized with 0.15 M NaCl ions. Simulation Parameters: The CHARMM36m force field [70] was employed for all protein, lipid, and ion parameters. Simulations were performed using GROMACS 2024.2 [71]. The system underwent energy minimization followed by a multi-step equilibration protocol involving NVT and NPT ensembles with position restraints on the protein heavy atoms, which were gradually released. Production simulations were run under the NPT ensemble at 303.15 K (using the v-rescale thermostat [72]) and 1.0 bar (using the C-rescale barostat [73] with semi-isotropic coupling). A 2 fs timestep was used, with LINCS algorithm [74] constraining bonds involving hydrogen atoms. Electrostatic interactions were calculated using the Particle Mesh Ewald (PME) method [75]). Trajectory Processing: Ten independent production replicas, each 2 $\mu$s in length, were generated. One replica exhibiting representative dynamics was selected for this study. The initial $\approx 776$ ns of this replica were discarded as extended equilibration, yielding a final analysis trajectory of $\approx 1.224$ $\mu$s. From this, 12,241 frames were sampled at a regular interval of 100 ps. All protein heavy-atom coordinates in these frames were then rigidly aligned to the heavy atoms of the first frame using the Kabsch algorithm [56] to remove global translation and rotation. Data Splitting and Preprocessing: The aligned coordinate dataset was split into a training set (90%, 11,017 frames) and a test set (10%, 1,224 frames) chronologically. Static topological information, including lists of backbone and sidechain atom indices based on a consistent re-indexing scheme, and the definitions of atom quadruplets for standard dihedral angles ($\phi,\psi,\chi_{1}-\chi_{5}$), was extracted once from the reference PDB structure. This information was stored in JSON format as detailed in Appendix B.
@@ F.2.1 Coordinate Accuracy
  • Local Distance Difference Test (lDDT) [76]: lDDT evaluates the preservation of local interatomic distances. For each atom, it considers all other atoms within a defined cutoff radius (e.g., 15 Å) in the reference (true) structure. It then calculates the fraction of these interatomic distances that are preserved in the predicted structure within certain tolerance thresholds (e.g., 0.5, 1, 2, and 4 Å). The final lDDT score is an average over these fractions and all residues/atoms. We report lDDTAll (all heavy atoms on backbone and sidechain) and lDDTBB (backbone heavy atoms). Scores range from 0 to 1, with 1 indicating perfect preservation.
  • Template Modeling score (TM-score) [77]: TM-score measures the global structural similarity between a predicted model and a reference structure. It is designed to be more sensitive to correct global topology and less sensitive to local errors than RMSD, and its value is normalized to be between 0 and 1, where 1 indicates a perfect match. A TM-score   0.5 generally indicates that the two proteins share a similar fold.
@@ F.2.3 Physical Plausibility
  • Average Steric Clash Counts: A steric clash is defined as a pair of non-bonded heavy atoms being closer than a specified distance cutoff. For our analysis, we used a cutoff of 2.1 Å. The clash score is the average number of such clashing pairs per generated structure. This metric was computed using BioPython [78] and SciPy’s [79] cKDTree for efficient neighbor searching.
@@ F.2.4 Conformational Landscape Sampling
[…]
@@ F.3 Training resources
The input Molecular Dynamics (MD) data was generated using GROMACS 2024.2 [71] on NVIDIA L40S GPUs. With a simulation rate of approximately 250 ns/day, the 2 micro seconds trajectory used in this study required about 8 days of computation. Following data generation, all machine learning models were implemented in PyTorch [58] and trained using the Adam optimizer [59] on NVIDIA L40S and H100 GPUs. The ChebNet autoencoder training required approximately 5,000 epochs at roughly 2 seconds per epoch. Decoder training varied by pooling strategy: blind pooling decoders trained for 3,000-6,000 epochs at approximately 3 seconds per epoch, while residue pooling decoders took about 24 seconds per epoch for a similar number of epochs. Diffusion model training was the most computationally intensive: sequential and blind pooling models trained for around 10,000 epochs, and residue pooling models for up to 150,000 epochs, with each epoch averaging approximately 1 second on L40S GPUs. The machine learning aspects presented in this paper required approximately 15,000 GPU hours, with total experimentation including preliminary setups amounting to roughly 30,000 GPU hours. This substantial GPU usage translates to significant energy consumption and associated carbon emissions; while precise quantification depends on factors like data center Power Usage Effectiveness (PUE) and local energy grid carbon intensity (and was not performed for this study), we acknowledge this environmental cost. Further details on computational resources, including specific MD simulation parameters and machine learning training runtimes for hyperparameter sweeps, are provided in the Extended Technical Appendix (Appendix I). The overall three-phase training and generation workflow (encoder pre-training, decoder training, and diffusion model training) is detailed in Appendix A (Algorithm 1).
@@ Appendix H Residue Pooling Inference Details
For the residue pooling model visualizations (Fig. 2c and Fig. 4c in the main text), samples were aggregated from 10 distinct model checkpoints saved at different stages of the diffusion model’s training. This multi-epoch sampling strategy, inspired by moving average techniques [80], was employed to provide a more stable and representative visualization of the learned conformational space, averaging out potential epoch-specific biases.
@@ I.2 Estimated CO2-equivalent emissions
Following [81], the per-GPU-hour emissions are:
@@
Board powers were taken from the NVIDIA L40S datasheet ($P=350$ W) [82] and the H100 SXM-5 specification ($P=700$ W) [83]. We adopt a conservative mean utilisation $u=0.8$, and a modern colocation PUE of 1.30, which is below the current global fleet average of 1.56 but well within the range observed for new facilities [84]. For the electricity mix we give two reference cases: (i) the 2022 global mean ($\mathrm{CI}=0.436$ kg CO2e kWh-1) derived from Ember’s Global Electricity Review [85], and (ii) the Swiss consumption-based factor of 0.112 kg CO2e kWh-1 for the same year, which captures the low-carbon hydro–nuclear mix and net imports [86].
@@ Results.
Table 4 summarises six scenarios. The ”all-L40S” and ”all-H100” scenarios are hypotheticals where the entire 30,000 GPU-hours were run on a single GPU type. More realistically, a mix of GPUs was used. For an illustrative scenario, we consider a split where 70% of the total operational hours (21,000 hours) were logged on L40S GPUs and 30% (9,000 hours) on H100 GPUs. The energy consumption for this specific split is calculated as the sum of energy consumed by each GPU type for their respective hours. Even the worst-case (all-H100 on the global-average grid) remains on the order of a single round-trip transatlantic flight for two passengers, while execution on a Swiss-like grid cuts emissions by $\sim$4$\times$.
@@ I.6 Latent Diffusion Models: Configuration and Training Overview
The Denoising Diffusion Probabilistic Models (DDPMs) [47], as outlined in Appendix D, were implemented to operate on the pooled latent embeddings $\mathbf{h}_{0}$ derived from each specific pooling strategy (with dimensions $d_{p}=100$ for Blind, $d_{p}=135$ for Sequential Backbone, $d_{p}=108$ for Sequential Sidechain, and $d_{p}=1092$ for Residue, as detailed in Section I.5). Prior to input, these $\mathbf{h}_{0}$ embeddings were normalized (zero mean, unit variance) based on statistics from the training set.
@@ I.9 Conditioning Mechanism Tuning
            self.q=nn.Linear(q_dim,att_dim,bias=False)
            self.k=nn.Linear(kv_dim,att_dim,bias=False)
            self.v=nn.Linear(kv_dim,att_dim,bias=False)
@@
Table 17: Ablation Study Results for Decoder Conditioning Mechanism. Performance is compared based on minimum test loss. ‘$X_{\text{ref}}$ Cond.‘ refers to conditioning on raw Cartesian coordinates, while ‘$Z_{\text{ref}}$ Cond.‘ refers to conditioning on latent reference embeddings. “1level” and “2level” refer to specific variants of the Residue Pooling architecture, with “2level” involving an initial global blind pooling stage prior to residue-specific processing.
@@
The results presented in Table 17 consistently show that conditioning on the latent reference embeddings ($Z_{\text{ref}}$) leads to substantially lower minimum test losses compared to conditioning on raw Cartesian coordinates ($X_{\text{ref}}$). This performance advantage holds across both Blind and Residue pooling strategies (including its “1level” and “2level” variants) and is evident whether Cross-Attention is used or not. For instance, with Blind Pooling and CA off, $Z_{\text{ref}}$ conditioning achieved a $\sim$31.4lower loss. This effect was even more pronounced for Residue Pooling (CA On), where an $\sim$80.2improvement was observed. Even for the more complex Residue (2level) architecture, $Z_{\text{ref}}$ conditioning provided a notable $\sim$24.3reduction in loss.
@@ I.10 Discussion on Metric Variability and Standard Deviations
[…]

### v4

@@ A.1 Notation
  • NN: Total heavy atoms per structure.
@@
  • $\mathcal{R}=\{R_{1},...,R_{N_{\text{res}}}\}$: Partition of atoms into residues.
@@ A.2 Overall Training and Generation Workflow
9:Train $\text{Encoder}_{\theta}$ on GraphDatasetGraphDataset using coordinate reconstruction loss $\mathcal{L}_{\text{HNO}}$ (Eq. in Appendix E).
@@
18:  Initialize $\text{Decoder}_{\phi}\leftarrow\text{BlindPoolingDecoder}(\dots)$.
@@
28:  Initialize $\text{Decoder}_{\phi}\leftarrow\text{ResidueBasedDecoder}(\dots)$.
@@
37:Train $\epsilon_{\theta}$ on DiffusionInputDatasetDiffusionInputDataset using $\mathcal{L}_{\text{diffusion}}$ (Eq. 2).
@@ Appendix B Input Data Processing and Representation
  • Zero-Based Contiguous Atom Indexing: A new, zero-based, and contiguous indexing scheme ($0,1,\dots,N-1$) is created for all $N$ heavy atoms in the protein system. This re-indexing maps original PDB atom identifiers to a consistent integer range, essential for constructing graph inputs and feature matrices for the neural network.
  • Residue Definitions: Residues are also re-indexed contiguously (e.g., $0,\dots,N_{\text{res}}-1$). For each re-indexed residue, the file stores:
@@
  • Dihedral Angle Definitions (Atom Quadruplets): This is a critical component for calculating dihedral-based losses and analyses. For each residue, the file stores definitions for all applicable standard backbone angles ($\phi,\psi$) and sidechain angles ($\chi_{1},\chi_{2},\chi_{3},\chi_{4},\chi_{5}$). Each dihedral angle is defined by an ordered quadruplet of four atom indices. Crucially, these atom indices adhere to the new, zero-based, contiguous indexing scheme. For example, a $\phi$ angle for a specific residue would be defined by four specific integer indices from the $0\dots N-1$ range. This allows for unambiguous calculation of any dihedral angle directly from a set of $N$ atomic coordinates, whether they are ground truth or model-predicted. The definitions also account for residue types where certain $\chi$ angles are not present (e.g., Alanine has no $\chi$ angles, Glycine has no sidechain).
@@ Appendix D Latent Diffusion Model Details
This section details the Denoising Diffusion Probabilistic Model (DDPM) [Ho et al., 2020] utilized in our framework. The DDPM operates on the pooled latent embeddings $\mathbf{h}_{0}\in\mathbb{R}^{d_{p}}$ (where $d_{p}$ is the dimension of the pooled latent space, dependent on the pooling strategy) obtained from the encoder and pooling stages. Specific architectures and hyperparameters for the denoising network $\epsilon_{\theta}$ are discussed in the Extended Technical Appendix (Section I).
@@ F.1 Dataset and MD Simulation Protocol
[…]
@@ F.2.1 Coordinate Accuracy
  • Local Distance Difference Test (lDDT) [Mariani et al., 2013]: lDDT evaluates the preservation of local interatomic distances. For each atom, it considers all other atoms within a defined cutoff radius (e.g., 15 Å) in the reference (true) structure. It then calculates the fraction of these interatomic distances that are preserved in the predicted structure within certain tolerance thresholds (e.g., 0.5, 1, 2, and 4 Å). The final lDDT score is an average over these fractions and all residues/atoms. We report lDDTAll (all heavy atoms on backbone and sidechain) and lDDTBB (backbone heavy atoms). Scores range from 0 to 1, with 1 indicating perfect preservation.
  • Template Modeling score (TM-score) [Zhang and Skolnick, 2005]: TM-score measures the global structural similarity between a predicted model and a reference structure. It is designed to be more sensitive to correct global topology and less sensitive to local errors than RMSD, and its value is normalized to be between 0 and 1, where 1 indicates a perfect match. A TM-score   0.5 generally indicates that the two proteins share a similar fold.
@@ F.2.3 Physical Plausibility
  • Average Steric Clash Counts: A steric clash is defined as a pair of non-bonded heavy atoms being closer than a specified distance cutoff. For our analysis, we used a cutoff of 2.1 Å. The clash score is the average number of such clashing pairs per generated structure. This metric was computed using BioPython [Cock et al., 2009] and SciPy’s [Virtanen et al., 2020] cKDTree for efficient neighbor searching.
@@ F.2.4 Conformational Landscape Sampling
  • A100 Activation Index Value: The A100 value is a collective variable developed by Ibrahim et al. [Ibrahim et al., 2019] to quantify the activation state of Class A G-protein-coupled receptors (GPCRs). It is a linear combination of five specific interhelical $\alpha-C$$\alpha$ distances that are known to change upon GPCR activation. The formula is given by:
@@ F.3 Training resources
The input Molecular Dynamics (MD) data was generated using GROMACS 2024.2 [Abraham et al., 2015] on NVIDIA L40S GPUs. With a simulation rate of approximately 250 ns/day, the 2 micro seconds trajectory used in this study required about 8 days of computation. Following data generation, all machine learning models were implemented in PyTorch [Paszke, 2019] and trained using the Adam optimizer [Kingma and Ba, 2014] on NVIDIA L40S and H100 GPUs. The ChebNet autoencoder training required approximately 5,000 epochs at roughly 2 seconds per epoch. Decoder training varied by pooling strategy: blind pooling decoders trained for 3,000-6,000 epochs at approximately 3 seconds per epoch, while residue pooling decoders took about 24 seconds per epoch for a similar number of epochs. Diffusion model training was the most computationally intensive: sequential and blind pooling models trained for around 10,000 epochs, and residue pooling models for up to 150,000 epochs, with each epoch averaging approximately 1 second on L40S GPUs. The machine learning aspects presented in this paper required approximately 15,000 GPU hours, with total experimentation including preliminary setups amounting to roughly 30,000 GPU hours. This substantial GPU usage translates to significant energy consumption and associated carbon emissions; while precise quantification depends on factors like data center Power Usage Effectiveness (PUE) and local energy grid carbon intensity (and was not performed for this study), we acknowledge this environmental cost. Further details on computational resources, including specific MD simulation parameters and machine learning training runtimes for hyperparameter sweeps, are provided in the Extended Technical Appendix (Appendix I). The overall three-phase training and generation workflow (encoder pre-training, decoder training, and diffusion model training) is detailed in Appendix A (Algorithm 1).
@@ Appendix H Residue Pooling Inference Details
For the residue pooling model visualizations (Fig. 2c and Fig. 4c in the main text), samples were aggregated from 10 distinct model checkpoints saved at different stages of the diffusion model’s training. This multi-epoch sampling strategy, inspired by moving average techniques [Qian et al., 2024], was employed to provide a more stable and representative visualization of the learned conformational space, averaging out potential epoch-specific biases.
@@ I.2 Estimated CO2-equivalent emissions
Following Henderson et al. [2020], the per-GPU-hour emissions are:
@@
Board powers were taken from the NVIDIA L40S datasheet ($P=350$ W) l [40] and the H100 SXM-5 specification ($P=700$ W) h [10]. We adopt a conservative mean utilisation $u=0.8$, and a modern colocation PUE of 1.30, which is below the current global fleet average of 1.56 but well within the range observed for new facilities PUE . For the electricity mix we give two reference cases: (i) the 2022 global mean ($\mathrm{CI}=0.436$ kg CO2e kWh-1) derived from Ember’s Global Electricity Review emb , and (ii) the Swiss consumption-based factor of 0.112 kg CO2e kWh-1 for the same year, which captures the low-carbon hydro–nuclear mix and net imports ali .
@@ Results.
[…]
@@ I.6 Latent Diffusion Models: Configuration and Training Overview
The Denoising Diffusion Probabilistic Models (DDPMs) [Ho et al., 2020], as outlined in Appendix D, were implemented to operate on the pooled latent embeddings $\mathbf{h}_{0}$ derived from each specific pooling strategy (with dimensions $d_{p}=100$ for Blind, $d_{p}=135$ for Sequential Backbone, $d_{p}=108$ for Sequential Sidechain, and $d_{p}=1092$ for Residue, as detailed in Section I.5). Prior to input, these $\mathbf{h}_{0}$ embeddings were normalized (zero mean, unit variance) based on statistics from the training set.
@@ I.9 Conditioning Mechanism Tuning
            self.q=$n$.Linear(q_dim,att_dim,bias=False)
            self.k=$n$.Linear(kv_dim,att_dim,bias=False)
            self.v=$n$.Linear(kv_dim,att_dim,bias=False)
@@
Table 17: Ablation Study Results for Decoder Conditioning Mechanism. Performance is compared based on minimum test loss. $X_{\text{ref}}$ Cond.‘ refers to conditioning on raw Cartesian coordinates, while $Z_{\text{ref}}$ Cond.‘ refers to conditioning on latent reference embeddings. “1level” and “2level” refer to specific variants of the Residue Pooling architecture, with “2level” involving an initial global blind pooling stage prior to residue-specific processing.
@@
The results presented in Table 17 consistently show that conditioning on the latent reference embeddings ($Z_{\text{ref}}$) leads to substantially lower minimum test losses compared to conditioning on raw Cartesian coordinates ($X_{\text{ref}}$). This performance advantage holds across both Blind and Residue pooling strategies (including its “1level” and “2level” variants) and is evident whether Cross-Attention is used or not. For instance, with Blind Pooling and CA off, $Z_{\text{ref}}$ conditioning achieved a $\sim$31.4% lower loss. This effect was even more pronounced for Residue Pooling (CA On), where an $\sim$80.2% improvement was observed. Even for the more complex Residue (2level) architecture, $Z_{\text{ref}}$ conditioning provided a notable $\sim$24.3% reduction in loss.
@@ I.10 Discussion on Metric Variability and Standard Deviations
In summary, while mean values are prioritized in the main tables for direct comparison of central tendencies, the analysis $\sumJSD fo$r dihedral distributions, PCA of latent spaces, A100 activation profiles, and the contextualized standard deviations discussed here collectively provide a comprehensive assessment of both the fidelity and the diversity of the conformational ensembles generated by LD-FPG.